# LUT-LLM: Efficient Large Language Model Inference with Memory-based Computations on FPGAs

### v1

@@ 3.2. Roofline Analysis and Comparison
Compared with the conventional roofline analyses of transformer acceleration (williams2009roofline; chen2024understanding) with arithmetic-based computations, we observe two properties of the naive memory-based computation for vector quantized model. First, memory-based compute can have higher throughput than arithmetic-based compute throughput with a longer sequence in the prefill stage, leading to at most $1.7\times$ performance boost compared to arithmetic-based compute. This is because memory-based compute utilizes memory access to compute dot products, increasing the total available compute resources. The throughput converges to the compute bound as the sequence length increases since the computation is dominated by attention layers. Second, the memory-based compute for Qwen 3 model needs to load lookup tables and codebooks that are $\mathbf{16\times}$ larger than FP16 weights. This results in significantly lower operational intensity and throughput in decode and short-context prefill compared to arithmetic-based compute. Such behavior is unfavorable for tasks involving short inputs or long outputs (e.g., short question-answering (rajpurkar2016squad) and creative writing (creative-writing-bench-v3)).
Thus, we investigate other quantization schemes for a more efficient alternative. Figure 6 compares the normalized throughput of various vector quantization methods with conventional scalar quantization on FPGA accelerators (hong2022dfx; he2025intar; chen2024allo), using $G=512$ and $c_{w}=16$ for weight quantization. Weight-only quantization boosts decoding throughput by reducing off-chip memory access, but its prefill performance remains constrained by costly floating-point operations, and decoding is further limited by restricted on-chip memory ports. In contrast, activation-weight co-quantization realizes the highest throughput across both stages with various input lengths. This improvement arises from three factors: (1) multiple weight vectors map to the same centroid vector, reducing the total number of entries required in the lookup tables and easing the off-chip memory bandwidth demand; (2) co-quantization lowers memory port requirements compared to weight-only or activation-only quantization to sustain high parallelism; and (3) INT8 table lookup and accumulation reduces on-chip memory pressure and eliminates frequent floating-point operations inherent in the dequantization of W4A8 scheme.
@@
Figure 6. Normalized throughputs of the prefill and decode stage vs. sequence length for Qwen 3 1.7B on AMD V80 FPGA with various quantization schemes. Activation-weight VQ can achieve superior performance than other schemes in both prefill and decode.
4. LUT-LLM Architecture
4.1. Architecture Overview
@@
Based on the performance analysis, we design the LUT-LLM accelerator with activation-weight co-quantization applied. Since not all operations can be executed through table lookup and the language model involves mixed precisions and computation orders, we carefully orchestrate the execution of operations and allocate on-chip resources to ensure its feasibility on the target FPGA. LUT-LLM integrates a LUTLinear engine with a global on-chip buffer, a dataflow attention engine, and two special function units (SFUs) for non-linear operations (SwiGLU and LayerNorm). The LUTLinear engine performs all linear projections and routes the outputs. The dataflow attention engine employs two FP32 GEMM/GEMV engines to compute $QK^{T}$ and $\text{softmax}(QK^{T})V$ in a pipelined manner, with rotational embedding and softmax applied in between. It supports both GEMM and GEMV to sustain high throughput in both the prefill and decode stages within a single design. The SFUs process the output embeddings from the LUTLinear engine according to the computations defined in the target model. In contrast to prior accelerators that need seperate designs for the prefill and decode stages, LUT-LLM is a unified design for both stages with overlapped KV-cache communication, computation, and data loading.
@@
Figure 7. The overall architecture of LUT-LLM, including a LUTLinear engine with global buffer, a dataflow attention engine, and special functions (SwiGLU, LayerNorm) with pipelined operations.
@@
Figure 8. The LUTLinear Engine (b) performs all linear projections with bandwidth-aware parallel centroid search units (BPCSU) and 2D lookup table prefix-sum (2D LUT PSum) engines, which process multiple input vectors on the hidden dimension. The accumulator and dequantizer is used to aggregate and convert results to FP32. (a) BPCSU uses distance PEs (dPE) and a reduction tree to the find nearest centroids, with parallelism determined by the allocated memory bandwidth. (c) The 2D LUT PSum engine retrieves centroid indices, gathers rows from the lookup tables, expands them with multiplexers, and performs SIMD addition before cascading to the next engine.
@@
The LUTLinear engine (Figure 8(b)) consists of two primary components: bandwidth-aware parallel centroid search units (BPCSU) and a 2D lookup table prefix-sum engines (2D LUT PSum Engine). BPCSU performs a nearest-neighbor search to identify the closest centroid for the given input vectors, while the 2D LUT PSum engine conducts a table lookup and accumulation to realize matrix multiplication. Each BPCSU is paired with a 2D LUT PSum engine, which receives the indices of the closest centroids. The 2D LUT PSum engines are connected in a serial chain to cascade and accumulate partial sums. Multiple BPCSUs and 2D LUT PSum engines enable parallel processing with multiple codebooks along the hidden dimension. The accumulated results are stored in the accumulator along with the output buffer. Before distributing outputs to other modules, the dequantizer converts each value in the output buffer to FP32 with the specified per-tensor scaling and shifting factors.
4.2. Bandwidth-aware Parallel Centroid Search
As shown in Figure 4, during inference, the first step is finding the closest centroids for the incoming input vectors. For hardware efficiency, the distance metric that is often used is the Chebyshev distance (cantrell2000modern), which calculates the distance between two vectors $\vec{a}$ and $\vec{b}$ by
(11) $L_{\infty}=\max_{i}(|a_{i}-b_{i}|)$
@@ 4.2. Bandwidth-aware Parallel Centroid Search
where $a_{i}$ and $b_{i}$ are elements in $\vec{a}$ and $\vec{b}$. To facilitate distance computation and comparison, prior works (li2025lut; wei2025t) instantiate distance PEs (dPEs) for every centroid in the codebook. Since codebooks are varied across hidden dimensions and layers, every dPE needs to load new centroids. dPEs are connected in a single-chain of pipeline, where the last dPE in the chain produces the index of the closest centroid. Although having low overhead when a long sequence of input vectors pipelined, this design incurs high setup latency if the sequence is short, e.g., in the decode stage. This latency is often higher than the off-chip memory access latency of lookup tables for LLMs. One potential alternative is to launch all dPEs to compute distance with all centroids in parallel, and apply a binary reduction tree to find the closest centroid. However, this significantly increases fanout and introduces $c_{a}-1$ comparators, which escalate routing difficulty, negatively impact timing, and spend more power on driving excessive resources.
In LUT-LLM, we propose the bandwidth-aware parallel centroid search unit (BPCSU), where we design the structure of data communication among dPEs to match the bandwidth of data loading in the 2D LUT PSum engine. We found that a complete reduction tree is often an overkill: we can support longer pipeline depth while the latency can still completely overlap with the off-chip memory access time. As seen in Figure 8(a), BPCSU arranges dPEs into multiple pipeline chains. The input vector is broadcasted to all chains for parallel distance computations. Each pipeline chain produces the minimum distance and index locally, which will be collected and compared through a small reduction tree. Consequently, BPCSUs can have a lower pipeline setup latency than data loading, with fewer resources instantiated for comparators than complete binary reduction trees. To maximize performance by overlapping centroid search and data loading, we have the following formulation: given $M$ as the output size, $G$ as the weight quantization group size, $c_{w},c_{a}$ are the weight and activation codebook sizes, and $C$ is the memory bandwidth per BPCSU (bits/cycle), the length of each dPE pipeline chain $l$ is
(12) $\arg\max_{l}((\frac{8c_{a}c_{w}M}{G}+\log(c_{w})M)/C\geq 32c_{a}/C+l+\log(c_{a}/l))$
where lookup tables are in INT8 and centroids are in FP32. $\log(c_{a}/l)$ is the depth of the reduction tree. We evaluate the maximum $l$ that satisfies this condition for every linear projection and select the minimum over them. For example, for configurations of Qwen 3 1.7B, $l=16$ for $c_{a}=64$ and $c_{w}=16$, indicating the instantiation of $16\times 4$ array of dPEs with 2-level reduction tree for each BPCSU.
4.3. Efficient 2D Table lookup Based Prefix-Sum
@@
Figure 9. (a) Partial sum with 1D lookup tables requires more memory capacity to store the tables and longer latency to access data for accumulation. (b) Partial sum in 2D LUT PSum engine needs smaller memory and fewer memory ports to have high parallelism in accumulation.
@@
After retrieving the closest centroid index, LUT-LLM will perform table lookups and accumulations to compute matrix multiplications. Naive memory-based computation accesses 1D lookup tables for every element in the output dimension. With a limited number of memory ports, e.g., 3 parallel access in Figure 9(a), the partial sum takes multiple cycles. To execute table lookup efficiently, we construct 2D lookup tables with a dedicated engine. Illustrated in Figure 9(b), given the incoming input centroid index received from BPCSUs, the 2D LUT PSum engine first accesses the row of 2D lookup tables with the corresponding index, and extracts to a LUT row register. Meanwhile, the centroid indices of the weight vectors are prefetched and stored in a register of length-$G$, where $G$ is the quantization group size. Then, the dot product result of each weight vector is retrieved and expanded by looking up the LUT row register using the weight centroid index. In the example, the engine will copy the value at position 1 of the LUT row register to position 0 in the output register for the first index in the weight centroid index register. Both the extraction of LUT row registers and data copying for expanded output registers finish in a single cycle with a memory partitioning and where each step in 2D table lookup is pipelined.
@@
Figure 8(c) illustrates the hardware architecture of the 2D LUT PSum engine. The 2D LUT buffer, which stores collections of lookup tables for each codebook, is instantiated utilzing a mixture of BRAM, URAM, and LUTRAM to balance memory resources on the target FPGA. Once the LUT row registers are extracted, their contents are forwarded to value copy multiplexers, which select entries according to the weight centroid index registers. To alleviate excessive fanout from the LUT row registers to the multiplexers, the registers are duplicated such that each copy drives only a subset of multiplexers. The expanded outputs are then accumulated with the results cascaded from the preceding 2D LUT PSum engine through a SIMD adder. The SIMD adders are realized with a combination of DSPs configured in SIMD mode or LUT-based adders to balance the computational resources on the FPGA.
4.4. Spatial-Temporal Hybrid Design
Figure 10. Compared to pure (a) dataflow or (b) sequential executions, (c) spatial-temporal hybrid execution can perfectly pipeline a nearest-neighbor search while saving memory resources for a parallel table lookup.
@@ 4.4. Spatial-Temporal Hybrid Design
To integrate the LUTLinear engine into a complete language model accelerator with high efficiency, it is essential to carefully orchestrate inter-operation communication. Conventional language model accelerators typically adopt either pure dataflow (chen2024allo; ye2025streamtensor) or sequential (zeng2024flightllm; hong2022dfx) execution strategies. While both exhibit a strong performance for arithmetic-intensive operations, they are suboptimal when the LUTLinear engine is introduced. For example, in a linear projection followed by attention computations, pure dataflow execution (Figure 10(a)) instantiates separate hardware modules for each operation and streams partial outputs through FIFOs. In the case of LUTLinear, however, generating partial outputs for streaming requires iterating over input vectors along the hidden dimension. This introduces two inefficiencies: (1) the same codebook must be reloaded multiple times, and (2) the centroid search cannot be pipelined, as the codebooks vary across the hidden dimension. These issues result in a significant stall latency in the linear layer. On the other hand, applying a purely sequential execution strategy (Figure 10(b)) requires the accelerator to allocate large buffers with high memory partitioning to compute and store attention scores. Consequently, the available memory capacity and port bandwidth for table lookups in LUTLinear are reduced. Since decoding and short-sequence prefilling are primarily constrained by memory access and the computation of linear layers, these architectural compromises significantly degrade overall performance.
@@
Inspired by prior accelerators that adopt hybrid execution patterns (he2025intar; zhuang2024ssr), we design LUT-LLM to employ a spatial-temporal hybrid execution strategy (Figure 10(c)). In this design, the LUTLinear engine executes sequentially across different operations, while its outputs are delivered to subsequent modules, i.e., attention, SwiGLU, and LayerNorm, via dataflow execution. This approach enables the LUTLinear engine to iterate over input vectors along the sequence dimension, thereby sustaining a perfectly pipelined nearest-neighbor search without codebook reloads. At the same time, the memory requirements for attention remain the same as pure dataflow execution, saving 14% on-chip resources for high-throughput table lookups in LUTLinear compared to pure sequential execution. This hybrid design balances between compute efficiency and memory utilization, establishing a scalable design principle for integrating memory-centric computation in LUT-LLM.
4.5. Other Design Features

### v2

@@ III-B Analysis and Comparison
We extend the performance model of vector-quantized linear layers to a complete transformer architecture following [9], with additional modeling for newly introduced computation blocks in modern LLMs (e.g., GQA, SwiGLU, etc.). Following prior work [57, 54, 55, 33], attentions and non-linear operations require floating point for accuracy. Figure 5 depicts the roofline analysis of deploying the Qwen 3 1.7B model on the AMD V80 FPGA with a naive memory-based computation (Act. VQ) like [30, 57] and arithmetic-based computation (FP16). The quantization configuration is: $c_{a}=64$, $v=2$, INT8 lookup tables, and FP32 codebooks. Compared with conventional roofline analyzes of transformer acceleration [58, 9] with arithmetic-based computations, we observe two properties of the naive memory-based computation for the vector quantized model. First, memory-based compute can have higher throughput than arithmetic-based compute throughput with a longer sequence in the prefill stage, leading to at most $1.7\times$ performance boost compared to arithmetic-based compute. This is because memory-based compute utilizes memory access to compute dot products, increasing the total available compute resources. The throughput converges to the compute bound as the sequence length increases, since the computation is dominated by attention layers. Second, naive memory-based computing needs to load lookup tables and codebooks that are $\mathbf{16\times}$ larger than the weights of FP16. This results in significantly lower operational intensity and throughput in decode and short-context prefill. Such behavior is unfavorable for tasks involving short inputs or long outputs (e.g., short question-answering [46] and creative writing [43]).
@@
Figure 5: Normalized throughputs of the prefill and decode stage for Qwen 3 1.7B on AMD V80 FPGA with various quantization schemes. Activation-weight VQ can achieve superior performance than other schemes in both prefill and decode. FP8/FP16 are not natively supported by DSPs on V80 and require online conversion to FP32 or custom compute logics [34, 63, 24].
We then investigate other quantization schemes for a more efficient alternative on FPGA such as W4A8, FP8, Weight VQ, and Activation + Weight VQ. Figure 5 compares the normalized throughput of these vector quantization methods with conventional scalar quantization on FPGA accelerators [24, 22, 10], using $G=512$ and $c_{w}=16$ for weight VQ. Weight VQ is employed with either arithmetic operations used in most GPU implementation as illustrated in Figure 2, or with memory-based operations by table lookups [17]. Weight VQ with arithmetic operations boosts decoding throughput by reducing off-chip memory access, but its prefill performance remains constrained by costly floating-point operations, and decoding is further limited by on-chip memory bandwidth. Weight VQ with table lookups, while mitigating these issues, requires large table storage, similar to activation VQ. In contrast, activation-weight co-quantization realizes the highest throughput across both stages than other scalar and vector quantization schemes. This improvement arises from three factors: (1) multiple weight vectors map to the same centroid vector, reducing the total number of entries required in the lookup tables and easing the off-chip memory bandwidth demand; (2) co-quantization has higher data reuse of lookup tables than weight-only or activation-only quantization due to repeating access to the same centroid, lowering memory port requirements to sustain high parallelism; and (3) INT8 table lookup and accumulation reduce on-chip memory pressure and eliminate FP dequantization operations requried by W4A8 scheme, ensuring high computational throughput on FPGAs.
IV LUT-LLM Architecture
IV-A Architecture Overview
@@
Based on the performance analysis, we design the LUT-LLM accelerator with activation-weight co-quantization applied. Since not all operations can be executed through table lookup and the language model involves mixed precisions and computation orders, we carefully orchestrate the execution of operations and allocate on-chip resources to ensure its feasibility on the target FPGA. LUT-LLM integrates a LUTLinear engine with a global on-chip buffer, a dataflow attention engine, and two special function units (SFUs) for non-linear operations (SwiGLU and LayerNorm). The LUTLinear engine performs all linear projections and routes the outputs. The dataflow attention engine employs two FP32 GEMM/GEMV engines to compute $QK^{T}$ and $\text{softmax}(QK^{T})V$ in a pipelined manner, with rotational embedding and softmax applied in between. It supports both GEMM and GEMV to sustain high throughput in both the prefill and decode stages within a single design. The SFUs handle nonlinear operations to process the output embeddings from the LUTLinear engine.
@@
Figure 6: The overall architecture of LUT-LLM, including a LUTLinear engine with global buffer, a dataflow attention engine, and special functions (SwiGLU, LayerNorm) with pipelined operations.
@@
Figure 7: The LUTLinear Engine (b) performs all linear projections with bandwidth-aware parallel centroid search units (BPCSU) and 2D lookup table prefix-sum (2D-PSum) engines. The accumulator and dequantizer is used to aggregate and convert results to FP32. (a) BPCSU uses distance PEs (dPE) and a reduction tree to find the nearest centroids, with parallelism determined by the allocated memory bandwidth. (c) 2D-PSum retrieves centroid indices, gathers rows from the lookup tables, expands them with multiplexers, and performs SIMD addition before cascading to the next engine.
@@
The LUTLinear engine (Figure 7(b)) consists of two primary components: bandwidth-aware parallel centroid search units (BPCSU) and a 2D lookup table prefix-sum engine (2D-PSum). BPCSU performs a nearest-neighbor search to identify the closest centroid for the given input vectors, while the 2D-PSum conducts a table lookup and accumulation to realize matrix multiplication. Each BPCSU is paired with a 2D-PSum, which receives the indices of the closest centroids. The 2D-PSum are connected in a serial chain to cascade and accumulate partial sums. Multiple BPCSUs and 2D-PSum enable parallel processing with multiple codebooks along the hidden dimension. The accumulated results are stored in the accumulator along with the output buffer. Before distributing outputs to other modules, the dequantizer converts each value in the output buffer to FP32 with the specified per-tensor scaling and shifting factors.
IV-B Bandwidth-aware Parallel Centroid Search
@@
As shown in Figure 4, during inference, the first step is finding the closest centroids for the incoming input vectors. For hardware efficiency, the distance metric that is often used is the Chebyshev distance [8], which calculates the distance between two vectors $\vec{a}$ and $\vec{b}$ by $L_{\infty}=\max_{i}(|a_{i}-b_{i}|)$, where $a_{i}$ and $b_{i}$ are elements in $\vec{a}$ and $\vec{b}$. To facilitate distance computation and comparison, previous work [30, 57] instantiated distance PEs (dPEs) for each centroid in the codebook. Since codebooks are varied across hidden dimensions and layers, every dPE needs to load new centroids. dPEs are connected in a single-chain of pipeline, where the last dPE in the chain produces the index of the closest centroid. Although having low overhead when a long sequence of input vectors is pipelined, this design incurs high setup latency if the sequence is short, e.g., in the decode stage. This latency is often higher than the off-chip memory access latency of lookup tables for LLMs. One potential alternative is to launch all dPEs to compute distance with all centroids in parallel, and apply a binary reduction tree to find the closest centroid. However, this significantly increases fanout and introduces $c_{a}-1$ comparators, which escalate routing difficulty, negatively impact timing, and consume more power on driving resources.
@@
Figure 8: (a) Partial sum with 1D lookup tables requires more memory capacity to store the tables and longer latency to access data for accumulation. (b) Partial sum in 2D-PSum needs smaller memory and fewer memory ports to have high parallelism in accumulation.
In LUT-LLM, we propose the bandwidth-aware parallel centroid search unit (BPCSU), where we design the structure of data communication among dPEs to match the bandwidth of data loading in the 2D-PSum. We found that a complete reduction tree is often an overkill: we can support longer pipeline depth while the latency can still completely overlap with the off-chip memory access time. As seen in Figure 7(a), BPCSU arranges dPEs into multiple pipeline chains. The input vector is broadcast to all chains for parallel distance computations. Each pipeline chain produces the minimum distance and index locally, which will be collected and compared through a small reduction tree. Consequently, BPCSUs can have a lower pipeline setup latency than data loading, with fewer resources instantiated for comparators than complete binary reduction trees. To maximize performance by overlapping centroid search and data loading, we have the following formulation: given $M$ as the output size, $G$ as the weight VQ group size, $c_{w},c_{a}$ are the weight and activation codebook sizes, and $C$ is the memory bandwidth per BPCSU (bits/cycle), the length of each dPE pipeline chain $l$ is
$\arg\max_{l}(\frac{8c_{a}c_{w}M}{GC}+\log(c_{w})\frac{M}{C})\geq 32\frac{c_{a}}{C}+l+\log(\frac{c_{a}}{l}))$ (9)
where lookup tables are in INT8 and centroids are in FP32. $\log(c_{a}/l)$ is the depth of the reduction tree. We evaluate the maximum $l$ that satisfies this condition for every linear projection and select the minimum over them. For example, for configurations of Qwen 3 1.7B, $l=16$ for $c_{a}=64$ and $c_{w}=16$, indicating the instantiation of a $16\times 4$ array of dPEs with a 2-level reduction tree for each BPCSU.
IV-C Efficient 2D Table lookup Based Prefix-Sum
@@
After retrieving the closest centroid index, LUT-LLM will perform table lookups and accumulations to compute matrix multiplications. Naive memory-based computation accesses 1D lookup tables for every element in the output dimension. With a limited number of memory ports (Figure 8(a)), the partial sum takes multiple cycles. To execute table lookup efficiently, we construct 2D lookup tables for the dot-product outputs and perform lookups with a dedicated engine. Element $(i,j)$ in each table is the dot product between centroid $i$ in activation codebook and centroid $j$ in weight codebook. Illustrated in Figure 8(b), given the incoming input centroid index received from BPCSUs, the 2D-PSum first accesses the row of 2D lookup tables with the corresponding index, and extracts to a LUT row register. Meanwhile, the centroid indices of the weight vectors are prefetched and stored in a register of length-$G$, where $G$ is the quantization group size. Then, the dot product result of each weight vector is retrieved and expanded by looking up the LUT row register using the weight centroid index. In the example, the engine will copy the value at position 1 of the LUT row register to position 0 in the output register for the first index in the weight centroid index register. Both the extraction of LUT row registers and data copying for expanded output registers finish in a single cycle with memory partitioning and pipelined 2D table lookup.
@@
Figure 7(c) illustrates the hardware architecture of the 2D-PSum. The 2D LUT buffer, which stores collections of lookup tables for each codebook, is instantiated utilizing a mixture of BRAM, URAM, and LUTRAM to balance memory resources on the target FPGA. Once the LUT row registers are extracted, their contents are forwarded to value copy multiplexers, which select entries according to the weight centroid index registers. To alleviate excessive fanout from the LUT row registers to the multiplexers, the registers are duplicated so that each copy drives only a subset of multiplexers. The expanded outputs are then accumulated with the results cascaded from the preceding 2D-PSum through a SIMD adder.
IV-D Spatial-Temporal Hybrid Design
Figure 9: (a) Dataflow execution: centroid search cannot be pipelined across tokens, and repeated codebook reloading induces pipeline stalls. (b) Sequential execution: excessive memory resources are allocated to the attention block, which degrades linear projection throughput. (c) Hybrid execution: enables full token-level pipelining while minimizing memory footprint, allowing parallel table lookup and preserving high projection throughput.
@@ IV-D Spatial-Temporal Hybrid Design
To integrate the LUTLinear engine into a complete language model accelerator with high efficiency, it is essential to carefully orchestrate inter-operation communication. Conventional language model accelerators typically adopt either pure dataflow [10, 61] or sequential [63, 24] execution strategies. While both exhibit strong performance for arithmetic-intensive operations, they are suboptimal when the LUTLinear engine is introduced. For example, in a linear projection followed by attention computations, pure dataflow execution (Figure 9(a)) instantiates separate hardware modules for each operation and streams partial outputs through FIFOs. However, in the case of LUTLinear, generating partial outputs for streaming requires iterating over input vectors along the hidden dimension. This introduces two inefficiencies: (1) the same codebook must be reloaded multiple times, and (2) the centroid search cannot be pipelined, as the codebooks vary across the hidden dimension. These issues result in a significant stall latency in the linear layer. On the other hand, applying a purely sequential execution strategy (Figure 9(b)) requires the accelerator to allocate large buffers with high memory partitioning to compute and store attention scores. Consequently, the available memory capacity and port bandwidth for table lookups in LUTLinear are reduced, constraining the decoding and short-context prefilling speed since they are memory-bound.
@@
Inspired by previous accelerators [22, 65], we design LUT-LLM to employ a spatial-temporal hybrid execution strategy (Figure 9(c)). The LUTLinear engine executes sequentially across different operations, while its outputs are delivered to subsequent modules, i.e., attention, SwiGLU, and LayerNorm, via dataflow execution. The data transfer logic is coarse-grain reconfigured similar to [22] by execution control signal to the accmulator (Figure 7(b)). This approach enables the LUTLinear engine to iterate over input vectors along the sequence dimension, sustaining a perfectly pipelined centroid search without codebook reloads. At the same time, the memory requirements for attention remain the same as dataflow execution, saving 14% on-chip resources for high-throughput table lookups used in LUTLinear than sequential execution.
IV-E Other Design Features